A school contains 357 boys and 323 girls
If a student is chosen at random, what is the probability of choosing a boy and a girl

Found 4 solutions by josgarithmetic, math_tutor2020, ikleyn, Edwin McCravy:
Answer by josgarithmetic(39792)

Something is wrong in the question. By the way written, "a student " should mean exactly 1 student. This one student to be both a boy and a girl at the same time seems uncommon.

Answer by math_tutor2020(3835)


P(boy) = probability of a boy
P(girl) = probability of a girl

P(boy) = (number of boys)/(number of students)
P(girl) = (number of girls)/(number of students)

Example:
A very small school has 10 boys and 12 girls
P(boy) = 10/(10+12) = 10/22 = 5/11
P(girl) = 12/(10+12) = 12/22 = 6/11
These results aren't the final answers to your particular problem; however, you can use this example to answer your question.

Answer by ikleyn(53751)

.

In mathematical problems,  written by professionals,  they do not use such form as it is used in this post.

Professional writers take a special care in order for wording be smooth and a reader would not stumble over unsuccessful phrases.

Otherwise,  it is seen from the distance of thousand miles,  that this composition
is created by an amateur on his knees in a garage,  and it's not funny - it's sad,  instead.


Math problems should  TEACH and  INSPIRE - - - it is their major purpose - not to sow disappointment.



Answer by Edwin McCravy(20077)

If the problem had been this way:

A school contains 357 boys and 323 girls.
If 2 students are chosen at random, what is the probability of choosing 
a boy and a girl.

The numerator of the probability would be the number of successful ways:
Choose a boy 357 ways, then choose a girl 357 ways.  323x357 = 115311 
successful ways.

The denominator of the probability would be the number of ways any two 
students could be chosen.  That's C(357+323,2) = C(680,2) = 230869

So the probability would be 115311/230860 = 969/1940 = 0.4994845361, very 
close to 0.5.

Edwin